Jack’s age is 5 less than three times Rebecca’s age. The sum of their ages is 75. What is the age of each of them??
Answer by ikleyn(52786)

.

Let x be the Rebecca's age.

Then Jack's age is  3x-5.


From the condition, you have this equation for the sum of ages

    x + (3x-5) = 75.


Simplify and solve

    x + 3x = 75 + 5

     4x    =    80

      x    =    80/4 = 20.


ANSWER.  Rebecca is 20 years old.   Jack is 3*20-5 = 55 years old.

Solved.